# Local Loss Optimization in the Infinite Width: Stable Parameterization of Predictive Coding Networks and Target Propagation

### v1

@@ 1 Introduction
Deep learning has achieved remarkable performance by building upon the backpropagation (BP) algorithm and developing architectures specialized for it [Rumelhart et al., 1986, LeCun et al., 1998, 2015]. BP, however, is not always a suitable method for more general objectives, such as biologically plausible computation [Lillicrap et al., 2020, Bredenberg et al., 2024] or efficient distributed computation [Amid et al., 2022]. A representative alternative is local loss optimization, a type of credit assignment problem, in which loss functions are defined layer-wise, and targets are set locally. The basic formulation involves performing regression on target signals at each layer to reduce the global error across the entire network: Predictive Coding networks, usually referred to as PC, generate their targets through the internal dynamics of inference [Whittington and Bogacz, 2017, Song et al., 2020, Salvatori et al., 2023], while Target Propagation (TP) generates them using feedback networks [Bengio, 2014, Lee et al., 2015, Ernoult et al., 2022].
@@
In many cases, the use of local losses requires additional hyperparameters (HPs) and their careful tuning, making the algorithm configuration significantly more complicated compared to that of BP. For example, PC requires not only the usual HPs, such as learning rate and initialization of weight parameters, but also those for the inference phase, such as the initialization of the state and the number of inference sequences. These HPs are primary considerations and have been reported as critical for ensuring stable training behavior [Pinchetti et al., 2024, Alonso et al., 2024, Rosenbaum, 2022]. A few analyses have succeeded in providing theoretical intuition for such local learning algorithms by introducing specific conditions or additional corrections that bridge them to classical optimization formulations [Song et al., 2020, Alonso et al., 2022, Meulemans et al., 2020]. However, such conditions are not always met in practice and may not be commonly shared across the entire family of methods. To develop local learning that is more easily manageable across a broader range of settings, it is promising to establish a theoretical foundation that enables the analysis of natural learning dynamics under fewer constraints.
@@
For standard BP, deep learning theory offers insights into the universal properties of learning [Bahri et al., 2020, Bartlett et al., 2021]. A key research focus in this area is understanding learning in the infinite-width limit, including studies on neural tangent kernel (NTK) and feature learning regimes [Jacot et al., 2018, Chizat et al., 2019, Mei et al., 2018, Bordelon and Pehlevan, 2022b]. In particular, Yang and Hu [2021] provided a unified perspective on the parameterizations that realize these learning regimes and proposed maximal update parameterization ($\mu$P) as a unique scaling of HPs, such as random initialization and learning rates, that achieves feature learning in the infinite-width limit. Building on this developing theoretical foundation, we expect to gain universal insight into local learning, which has not yet been systematically analyzed.
@@
In this work, we derive the $\mu$P for PC and TP and investigate hyperparameter transfer (the so-called $\mu$Transfer) across different widths. Our contributions are summarized as follows:
@@
  • Second, for a more general context involving multiple inference sequences, we consider the convergence of the inference phase. We find that, for deep linear networks, we can explicitly obtain the local targets and losses at the fixed point of the inference, which depend on inference step sizes (Theorem 4.2). Interestingly, it takes a similar form to the Gauss-Newton (GN) gradient, but it can reduce to the conventional first-order gradient descent (GD) depending on the parameterization and step sizes. We find that the eventual gradient is closer to GD for sufficiently wide neural networks under standard experimental settings with $\mu$P. We also confirm that a larger inference step size, identified through this analysis, enhances $\mu$Transfer of HPs.
@@ 2 Related Work
Local learning: Most research on local learning stems from the exploration of biologically plausible learning [Lillicrap et al., 2020], with PC and TP following this line. As deep learning has evolved, local learning has also begun to focus on large-scale networks, and some models have achieved performances close to those trained with BP [Ernoult et al., 2022, Ren et al., 2023]. Several algorithms are inherently structured to resemble the BP chain [Akrout et al., 2019] or to estimate first-order gradients [Scellier and Bengio, 2017]. In contrast, PC relies on an inference phase, and TP uses a feedback network, both of which are quite different from BP and seems to be fundamental designs for using local targets. However, their optimization properties are still not well understood. Alonso et al. [2022] proposed a modified PC as a proximal point algorithm (implicit SGD), though it requires additional corrections and adaptive rescaling [Alonso et al., 2024]. Innocenti et al. [2023] proposed an inference phase computed by a GN method, but it requires a quadratic approximation of the local loss around a special initialization. As discussed in the next section, bridging to such classical optimization requires strong conditions that may deviate significantly from the original purpose and algorithm [Rosenbaum, 2022, Meulemans et al., 2020].
Infinite width and $\mu$P: While the NTK regime guarantees the existence of learning dynamics in the infinite-width limit and its global convergence, it reduces to just a kernel method [Jacot et al., 2018, Lee et al., 2019]. To realize feature learning in the infinite-width limit, Yang and Hu [2021] proposed $\mu$P, which is a non-trivial scaling of HPs with respect to the width. From a theoretical perspective, this serves as a parameterization that enables the dynamics of feature learning, such as those described by the mean-field regime [Mei et al., 2018] or the dynamical mean-field theory [Bordelon and Pehlevan, 2022b]. For more applications, $\mu$P or its extension has been validated across various architectures [Yang et al., 2021, Vyas et al., 2023, Everett et al., 2024]. It covers not only the naive first-order gradient but also entry-wise adaptive optimizers such as Adam [Yang and Littwin, 2023] and second-order optimization methods like K-FAC [Ishikawa and Karakida, 2024]. There has been little previous work on the infinite-width analysis of local learning. Bordelon and Pehlevan [2022a] formulated (direct) feedback alignment and (supervised) Hebbian learning using dynamical mean-field theory, which are rather close to BP.
@@ 3 Preliminaries
In this section, we provide an overview of local learning and $\mu$P in an $L$-layer fully connected neural network:
@@
where $W_{l}\in\mathbb{R}^{M_{l}\times M_{l-1}}$ are weight matrices, $h_{l},u_{l}\in\mathbb{R}^{M_{l}\times N}$ are activations and $N$ is the number of samples, independent of the order of width $M$. We set the width of the hidden layers to $M_{l}=M$ for $(l=1,\dots,L-1)$ for simplicity. To keep the notation concise, for non-linear networks, we set $M_{L}=1$ ; however, we can easily generalize to $M_{L}=\Theta(1)$. The activation function $\phi(\cdot)$ is usually assumed to be differentiable and polynomially bounded for some theoretical reasons within the $\mu$P framework [Yang and Hu, 2021].
@@ 3.1.1 Predictive Coding
Predictive Coding updates both the states and weights to minimize the following free-energy function [Whittington and Bogacz, 2017, Song et al., 2020, Salvatori et al., 2023]:
@@
To distinguish the internal state from the forward signal propagation $u_{l}$, we denote this state as $v_{l}$. Although this algorithm was originally derived from the variational Bayes formulation, it has been extended beyond the scope of the original framework, aiming instead to develop inference computations that work more effectively in practice. PC is composed of two phases: an inference phase, in which the per-layer states $v_{l}$ are updated and a learning phase, in which weights $W_{l}$ are updated. Its update rule for the inference phase is given by
$v_{l,s+1}$ $=v_{l,s}-\frac{\partial\mathcal{F}}{\partial v_{l}}=v_{l,s}-\gamma_{l}e_{l,s}+\gamma_{l+1}\phi^{\prime}\left(v_{l,s}\right)\circ W_{l+1}^{\top}e_{l+1,s},$ (3)
where we define $e_{l,s}:=v_{l,s}-W_{l}\phi(v_{l-1,s})$ and $\circ$ denotes the Hadamard product. The update rule for the learning phase is given by
@@
Note that the inference time index $s$ and the parameter update index $t$ are distinct with $s$ resetting to 0 at each $t$. We usually omit the step size $\gamma_{l}$ in Eq. (4) in implementation. Generally, weights are updated after multiple inference steps, while incremental version of PC (iPC), which updates the weights after just a single inference, has also been proposed [Salvatori et al., 2024]. The internal state can be updated simultaneously across all layers or computed sequentially in a specified order. In the first part of the next section, we focus on the Sequential Inference (SI) method, where $e_{l,s}$ is computed sequentially by propagating from the output layer to the input layer. For more details on this difference, see Algorithm 1 in the Appendix.
Empirically, to improve the trainability, PC often relies on assumptions that are either rational or, at times, unrealistic. One such reasonable assumption is the initialization method for $v_{l,0}$, which is used to improve the convergence [Song et al., 2020, Alonso et al., 2022, Rosenbaum, 2022]:
@@
Generally, the gradient computation of PC does not match that of BP. However, under F-ini and SI, it reduces to BP by adopting the following rather technical assumption [Millidge et al., 2022b, Rosenbaum, 2022]:
@@
Under FPA, the inference is given by ${{e}}_{l,s+1}=(1-\gamma_{l}){{e}}_{l,s}+\gamma_{l+1}\phi^{\prime}\left(v_{l,0}\right)\circ W_{l+1}^{\top}{{e}}_{l+1,s}.$ By substituting F-ini, one can easily verify that this sequential inference computes $\nabla_{u_{l}}\mathcal{L}$.
@@ 3.1.2 Target Propagation
where $Q_{l}\in\mathbb{R}^{M_{l-1}\times M_{l}}$ are weight matrices. We also analyze the Difference Target Propagation (DTP), a variant of TP, whose definition is provided in the appendix. The feedback network is trained to minimize the following reconstruction loss:
@@
where $f_{l}(x)=\phi(W_{l}x)$. TP updates the weights $W_{l}$ to minimize the following local loss $\|e_{l}\|^{2}:=\|\hat{h}_{l}-h_{l}\|^{2}$. The gradient of this local loss provides the update rule for the learning phase as $W_{l,t+1}=W_{l,t}-\eta_{l}\phi^{\prime}(W_{l,t}{h}_{l-1})\circ e_{l}{h}_{l-1}^{\top}.$ For a so-called invertible network, TP computes the Gauss-Newton Target (GNT), i.e., $e_{l}^{\text{GNT}}=(\delta_{l}\delta_{l}^{\top}+\rho I)^{-1}\delta_{l}{e}_{L}$ where $\delta_{l}=\nabla_{u_{l}}u_{L}$ is the BP signal [Meulemans et al., 2020]  [^1] The final gradient $d\mathcal{F}/dW_{l}$ is equivalent to the special case of K-FAC [Martens and Grosse, 2015] where the preconditioners are applied only to the backward signals.. Note that the assumption of the invertible network is restrictive because the invertible network requires invertible activation functions, regular weight matrices, and the training tp converge to the solution of $g_{l}(\hat{{h}}_{l+1})=f_{l+1}^{-1}(\hat{{h}}_{l+1})=W_{l+1}^{-1}\phi^{-1}(\hat{{h}}_{l+1})$. For general networks, (D)TP does not necessarily lead to the GNT.
Remark on a connection between PC and TP. Some previous studies have argued that PC yields GNT-like solutions, and thus can be connected to TP [Alonso et al., 2022, Millidge et al., 2022a]. These works attempt to gain an intuitive insight from the fixed point equation for each layer:
@@
where $h_{l}^{*}$ means $\phi(v_{l}^{*})$. For $\gamma_{l+1}/\gamma_{l}\ll 1$, we approximate $h_{l}^{*}\approx W_{l+1}^{\dagger}h_{l+1}^{*}$. If we multiply this approximation across layers, the naive expectation is that $h_{l}^{*}\approx\prod_{i=l+1}^{L}W_{i}^{\dagger}e_{L}^{*}$, which corresponds to the GNT for linear networks. Thus, we can intuitively see that the PC may be linked to the GNT, although its exact connection requires careful limit operations across layers. Additionally, taking the limits $\gamma_{l+1}/\gamma_{l}\ll 1$ for all layers means the exponential decay of $\gamma_{l}$ with depth, rasing concerns regarding its practical relevance. For $\gamma_{l}=1$, Innocenti et al. [2024] has recently derived an explicit formulation of the free energy at the fixed point using an unfolding calculation of a hierarchical Gaussian model. This fomulation shows that the obtained gradient differs from that of the exact GNT, supporting the idea that the connection to GNT would be weak.
@@ 3.2 $\mu$P and Learning Regimes
The abc-parameterization scales the parameters by width, as follows [Yang and Hu, 2021]:
@@
where $\Theta(\cdot)$ denotes the order with respect to the width and $x=\Theta(M^{a})$ means $\sqrt{\|x\|^{2}/M}=\Theta(M^{a})$ for $x\in\mathbb{R}^{M}$. Yang and Hu [2021] characterized the condition where $r_{l}=0$ for all layers as the feature learning regime and the condition where $r_{l<L}>0$ and $r_{L}$ = 0 as the kernel regime. The parametrization that induces the feature learning regime is called $\mu$P, while the NTK parametrization corresponds to the kernel regime with $r_{l<L}=1/2$. The training dynamics are referred to as stable when they neither vanish nor explode as the network width increases; the parameterization that achieves this is called a stable parameterization. Yang and Hu [2021] demonstrated that the concrete evaluation of feature learning with $r_{l}=0$ reduces to the following two conditions:
@@
Note that the original derivation of the parametrization that satisfies the above conditions is based on the first (infinitesimal) one-step update of the parameters [Yang and Hu, 2021, Ishikawa and Karakida, 2024]. Our work also follows the same approach.
3.2.1 $\mu$P for Gauss-Newton Target
The following work has recently derived the $\mu$P scaling, including both first-order and second-order optimizations.
Proposition 3.3 (Ishikawa and Karakida [2024]).
Consider the first one-step update by the GNT: $W_{l,1}=W_{l,0}-\eta_{l}\phi^{\prime}(W_{l,t}{h}_{l-1})\circ(\delta_{l}\delta_{l}^{\top}+\rho I)^{-e_{B}}\delta_{l}\text{diag}({e}_{L})h_{l-1}^{\top}$. In the infinite-width limit, this update admits the $\mu$P for feature learning at
@@ 3.2.1 $\mu$P for Gauss-Newton Target
Table 1: Parameterization ($b_{l},c_{l}$). PC with $\bar{\gamma}_{L}=0$ reduces to SGD’s $\mu$P, while one with $\bar{\gamma}_{L}=-1$ reduces to GNT’s. TP has the distinctive property of $b_{L}=1/2$.
Layer SP SGD GNT PC TP
@@
More precisely, we can also allow $b_{L}\geq 1$ for the feature learning regime. However such initialization eventually reduces to the case of $b_{L}=1$ in the next parameter update. Thus, we effectively summarize it as $b_{L}=1$. The scaling of $b_{L}=1$ implies that a smaller initialization is required compared to the standard parametrization (SP), which is PyTorch’s default, for sufficiently wide neural networks. It can also be immediately verified that we can set $a_{l}=0$ due to shift invariance without loss of generality. In Table 1, we summarize the $\mu$P from previous work and our results obtained in the following sections.

### v2

@@ 1 Introduction
Deep learning has achieved remarkable performance by building upon the backpropagation (BP) algorithm and developing architectures specialized for it (Rumelhart et al., 1986; LeCun et al., 1998; 2015). BP, however, is not always a suitable method for more general objectives, such as biologically plausible computation (Lillicrap et al., 2020; Bredenberg et al., 2024) or efficient distributed computation (Amid et al., 2022). A representative alternative is local loss optimization, a type of credit assignment problem, in which loss functions are defined layer-wise, and targets are set locally. The basic formulation involves performing regression on target signals at each layer to reduce the global error across the entire network: Predictive Coding networks, usually referred to as PC, generate their targets through the internal dynamics of inference (Whittington & Bogacz, 2017; Song et al., 2020; Salvatori et al., 2023), while Target Propagation (TP) generates them using feedback networks (Bengio, 2014; Lee et al., 2015; Ernoult et al., 2022).
@@
In many cases, the use of local losses requires additional hyperparameters (HPs) and their careful tuning, making the algorithm configuration significantly more complicated compared to that of BP. For example, PC requires not only the usual HPs, such as learning rate and initialization of weight parameters but also those for the inference phase, such as the initialization of the state and the number of inference sequences. These HPs are primary considerations and have been reported as critical for ensuring stable training behavior (Pinchetti et al., 2024; Alonso et al., 2024; Rosenbaum, 2022). A few analyses have succeeded in providing theoretical intuition for such local learning algorithms by introducing specific conditions or additional corrections that bridge them to classical optimization formulations (Song et al., 2020; Alonso et al., 2022; Meulemans et al., 2020). However, such conditions are not always met in practice and may not be commonly shared across the entire family of methods. To develop local learning that is more easily manageable across a broader range of settings, it is promising to establish a theoretical foundation that enables the analysis of natural learning dynamics under fewer constraints.
@@
For standard BP, deep learning theory offers insights into the universal properties of learning (Bahri et al., 2020; Bartlett et al., 2021). A key research focus in this area is understanding learning in the infinite-width limit, including studies on neural tangent kernel (NTK) and feature learning regimes (Jacot et al., 2018; Chizat et al., 2019; Mei et al., 2018; Bordelon & Pehlevan, 2022b). In particular, Yang & Hu (2021) provided a unified perspective on the parameterizations that realize these learning regimes and proposed maximal update parameterization ($\mu$P) as a unique scaling of HPs, such as random initialization and learning rates, that achieves feature learning in the infinite-width limit. Building on this developing theoretical foundation, we expect to gain universal insight into local learning, which has not yet been systematically analyzed.
@@
In this work, we derive the $\mu$P for PC and TP and investigate hyperparameter transfer (the so-called $\mu$Transfer) across different widths. Although $\mu$P for SGD has been previously derived, the $\mu$P depends on the specific training algorithm, making it necessary to derive $\mu$P for each local learning algorithm. Our contributions are summarized as follows:
@@
  • Second, for a more general context involving multiple inference sequences, we consider the convergence of the inference phase. We find that, for deep linear networks, we can explicitly obtain the local targets and losses at the fixed point of the inference, which depend on inference step sizes (Theorem 4.2). Interestingly, it takes a similar form to the Gauss-Newton (GN) gradient, but it can be reduced to the conventional first-order gradient descent (GD) depending on the parameterization and step sizes. We find that the eventual gradient is closer to GD for sufficiently wide neural networks under standard experimental settings with $\mu$P. We also confirm that a larger inference step size, identified through this analysis, enhances $\mu$Transfer of HPs.
@@ 2 Related Work
Local learning: Most research on local learning stems from the exploration of biologically plausible learning (Lillicrap et al., 2020), with PC and TP following this line. As deep learning has evolved, local learning has also begun to focus on large-scale networks, and some models have achieved performances close to those trained with BP (Ernoult et al., 2022; Ren et al., 2023). Several algorithms are inherently structured to resemble the BP chain (Akrout et al., 2019) or to estimate first-order gradients (Scellier & Bengio, 2017). In contrast, PC relies on an inference phase, which essentially infers the appropriate activation values for hidden layers, and TP uses a feedback network, both of which are quite different from BP and seem to be fundamental designs for using local targets. However, their optimization properties are still not well understood. Alonso et al. (2022) proposed a modified PC as a proximal point algorithm (implicit SGD), though it requires additional corrections and adaptive rescaling (Alonso et al., 2024). Innocenti et al. (2023) proposed an inference phase computed by a GN method, but it requires a quadratic approximation of the local loss around a special initialization. As discussed in the next section, bridging to such classical optimization requires strong conditions that may deviate significantly from the original purpose and algorithm (Rosenbaum, 2022; Meulemans et al., 2020).
Infinite width and $\mu$P: While the NTK regime guarantees the existence of learning dynamics in the infinite-width limit and its global convergence, it reduces to just a kernel method (Jacot et al., 2018; Lee et al., 2019). To realize feature learning in the infinite-width limit, Yang & Hu (2021) proposed $\mu$P, which is a non-trivial scaling of HPs with respect to the width. From a theoretical perspective, this serves as a parameterization that enables the dynamics of feature learning, such as those described by the mean-field regime (Mei et al., 2018) or the dynamical mean-field theory (Bordelon & Pehlevan, 2022b). For more applications, $\mu$P or its extension has been validated across various architectures (Yang et al., 2021; Vyas et al., 2023; Everett et al., 2024). It covers not only the naive first-order gradient but also entry-wise adaptive optimizers such as Adam (Yang & Littwin, 2023) and second-order optimization methods like K-FAC (Ishikawa & Karakida, 2024). There has been little previous work on the infinite-width analysis of local learning. Bordelon & Pehlevan (2022a) formulated (direct) feedback alignment and (supervised) Hebbian learning using dynamical mean-field theory, which are rather close to BP.
@@ 3 Preliminaries
In this section, we summarize local learning and $\mu$P in an $L$-layer fully connected neural network $f$:
@@
where $W_{l}\in\mathbb{R}^{M_{l}\times M_{l-1}}$ are weight matrices, $h_{l},u_{l}\in\mathbb{R}^{M_{l}\times N}$ are activations and $N$ is the number of data samples, independent of the order of width $M_{l}$. We set the width of the hidden layers to $M_{l}=M$ for $(l=1,\dots,L-1)$ for simplicity. To keep the notation concise, for non-linear networks, we set $M_{L}=1$; however, we can easily generalize to $M_{L}=\Theta(1)$. The activation function $\phi(\cdot)$ is usually assumed to be differentiable and polynomially bounded for some theoretical reasons within the $\mu$P framework (Yang & Hu, 2021).
@@ 3.1.1 Predictive Coding
Predictive Coding (PC) updates both the states and weights to minimize the following free-energy function (Whittington & Bogacz, 2017; Song et al., 2020; Salvatori et al., 2023):
@@
$\mathcal{L}$ denotes a loss function, and both mean squared error loss and cross-entropy loss are allowed in theory and experiments unless an explicit assumption is stated. To distinguish the internal state from the forward signal propagation $u_{l}$, we denote this state as $v_{l}$. Although this algorithm was originally derived from the variational Bayes formulation, it has been extended beyond the scope of the original framework, aiming instead to develop inference computations that work more effectively in practice. PC is composed of two phases: an inference phase, in which the per-layer states $v_{l}$ are updated and a learning phase, in which weights $W_{l}$ are updated. Its update rule for the inference phase is given by
$v_{l,s+1}$ $=v_{l,s}-\frac{\partial\mathcal{F}}{\partial v_{l}}=v_{l,s}-\gamma_{l}e_{l,s}+\gamma_{l+1}\phi^{\prime}\left(v_{l,s}\right)\circ W_{l+1}^{\top}e_{l+1,s}\quad(l<L),$ (3)
where we define $e_{l,s}:=v_{l,s}-W_{l}\phi(v_{l-1,s})$ and $\circ$ is the Hadamard product. From eq. (3), $\gamma_{l}$ can be regarded as a step size for the inference phase. The update rule for the learning phase is given by
@@
Note that the inference time index $s$ and the parameter update index $t$ are distinct with $s$ resetting to 0 at each $t$. We usually omit the step size $\gamma_{l}$ in Eq. (S.6) in implementation. Generally, weights are updated after multiple inference steps, while the incremental version of PC (iPC), which updates the weights after just a single inference, has also been proposed (Salvatori et al., 2024b). The internal state can be updated simultaneously across all layers or computed sequentially in a specified order. In the first part of the next section, we focus on the Sequential Inference (SI) method, where $e_{l,s}$ is computed sequentially by propagating from the output layer to the input layer. For more details on this difference, see Algorithm 1 in the Appendix.
Empirically, to improve trainability, PC often relies on assumptions that are either rational or, at times, unrealistic. One such reasonable assumption is the initialization method for $v_{l,0}$, which is used to improve the convergence (Song et al., 2020; Alonso et al., 2022; Rosenbaum, 2022):
@@
Generally, the gradient computation of PC does not match that of BP. However, under F-ini and SI, it reduces to BP by adopting the following rather technical assumption (Millidge et al., 2022b; Rosenbaum, 2022):
@@
Under FPA, the inference is given by ${{e}}_{l,s+1}=(1-\gamma_{l}){{e}}_{l,s}+\gamma_{l+1}\phi^{\prime}\left(v_{l,0}\right)\circ W_{l+1}^{\top}{{e}}_{l+1,s}.$ By substituting F-ini, one can easily verify that this sequential inference computes $\nabla_{u_{l}}\mathcal{L}$. In Section 4, we reveal that the following scaling of $\gamma_{L}$ with respect to the width $M$ plays a fundamental role in characterizing the feature learning of PC and a parameterization that enables stable learning even without such heuristic techniques:
$\gamma_{L}=\gamma\textquoteright/M^{\bar{\gamma}_{L}}$ (5)
[…]
@@ 3.1.2 Target Propagation
where $Q_{l}\in\mathbb{R}^{M_{l-1}\times M_{l}}$ are weight matrices, $\psi(\cdot)$ is an activation function of the feedback network. We also analyze the Difference Target Propagation (DTP), a variant of TP, whose definition is provided in the appendix. The feedback network is trained to minimize the following reconstruction loss:
@@
where $f_{l}(x)=\phi(W_{l}x)$. TP updates the weights $W_{l}$ to minimize the following local loss $\|e_{l}\|^{2}:=\|\hat{h}_{l}-h_{l}\|^{2}$. The gradient of this local loss provides the update rule for the learning phase as $W_{l,t+1}=W_{l,t}-\eta_{l}\phi^{\prime}(W_{l,t}{h}_{l-1})\circ e_{l}{h}_{l-1}^{\top}.$ For a so-called invertible network, TP computes the Gauss-Newton Target (GNT), i.e., $e_{l}^{\text{GNT}}=(\delta_{l}\delta_{l}^{\top}+\rho I)^{-1}\delta_{l}{e}_{L}$ where $\delta_{l}=\nabla_{u_{l}}u_{L}$ is the BP signal (Meulemans et al., 2020)  [^1] The final gradient $d\mathcal{F}/dW_{l}$ is equivalent to the special case of K-FAC (Martens & Grosse, 2015) where the preconditioners are applied only to the backward signals. and $e_{L}=y-h_{L}$ is the error vector. Note that the assumption of the invertible network is restrictive because the invertible network requires invertible activation functions, regular weight matrices, and the training tp converge to the solution of $g_{l}(\hat{{h}}_{l+1})=f_{l+1}^{-1}(\hat{{h}}_{l+1})=W_{l+1}^{-1}\phi^{-1}(\hat{{h}}_{l+1})$. For general networks, (D)TP does not necessarily lead to the GNT.
Remark on a connection between PC and TP. Some previous studies have argued that PC yields GNT-like solutions, and thus can be connected to TP (Alonso et al., 2022; Millidge et al., 2022a). These works attempt to gain an intuitive insight from the fixed point equation for each layer:
@@
where $h_{l}^{*}$ means $\phi(v_{l}^{*})$. For $\gamma_{l+1}/\gamma_{l}\ll 1$, we approximate $h_{l}^{*}\approx W_{l+1}^{\dagger}h_{l+1}^{*}$. If we multiply this approximation across layers, the naive expectation is that $h_{l}^{*}\approx\prod_{i=l+1}^{L}W_{i}^{\dagger}e_{L}^{*}$, which corresponds to the GNT for linear networks. Thus, we can intuitively see that the PC may be linked to the GNT, although its exact connection requires careful limit operations across layers. Additionally, taking the limits $\gamma_{l+1}/\gamma_{l}\ll 1$ for all layers means the exponential decay of $\gamma_{l}$ with depth, raising concerns regarding its practical relevance. For $\gamma_{l}=1$, Innocenti et al. (2024) has recently derived an explicit formulation of the free energy at the fixed point using an unfolding calculation of a hierarchical Gaussian model. This formulation shows that the obtained gradient differs from that of the exact GNT, supporting the idea that the connection to GNT would be weak.
@@ 3.2 $\mu$P and Learning Regimes
The abc-parameterization $\left\{a_{l},b_{l},c_{l}\right\}_{1\leq l\leq L}$ determines the scaling of weights and learning rates at initialization. It scales the parameters by width as follows (Yang & Hu, 2021):
@@
where $\Theta(\cdot)$ denotes the order with respect to the width and $x=\Theta(M^{a})$ means $\sqrt{\|x\|^{2}/M}=\Theta(M^{a})$ for $x\in\mathbb{R}^{M}$.The training dynamics and parameterization are referred to as stable when $u_{l,0}$ neither vanish nor explode as the network width increases and $\Delta h_{l,t}$ do not explode as the network width increases (Definition S.15). Yang & Hu (2021) introduced the following conditions and characterized $\mu$P as a unique stable abc-parameterization under them:
@@
These conditions imply $r_{l}=0$ for all layers and feature learning. In contrast to this feature learning regime, the previous work refers $r_{l<L}>0$ and $r_{L}$ = 0 as the kernel regime. The NTK parameterization corresponds to the kernel regime with $r_{l<L}=1/2$ Note that the original derivation of the parameterization that satisfies the above conditions is based on the first (infinitesimal) one-step update of the parameters (Yang & Hu, 2021; Ishikawa & Karakida, 2024) (see Section A.2). Our work also follows the same approach.
$\mu$P for Gauss-Newton Target: The following work has recently derived the $\mu$P scaling, including both first-order and second-order optimizations.
Proposition 3.3 (Ishikawa & Karakida (2024)).
Consider the first one-step update by the GNT: $W_{l,1}=W_{l,0}-\eta_{l}\phi^{\prime}(W_{l,t}{h}_{l-1})\circ(\delta_{l}\delta_{l}^{\top}+\rho I)^{-e_{B}}\delta_{l}\text{diag}({e}_{L})h_{l-1}^{\top}$ where $\delta_{l}=\nabla_{u_{l}}u_{L}$ and $e_{L}=y-h_{L}$. In the infinite-width limit, this update admits the $\mu$P for feature learning at
@@
Table 1: Parameterization for weight initialization scale: $b_{l}$ and learning rate scale: $c_{l}$. Predictive Coding (PC) with $\bar{\gamma}_{L}=0$ reduces to SGD’s $\mu$P, while one with $\bar{\gamma}_{L}=-1$ reduces to $\mu$P for Gauss Newton Target (GNT). TP (Target Propagation) has the distinctive property of $b_{L}=1/2$.
Layer SP (Default) SGD (2021) GNT (2024) PC (New) TP (New)
@@
More precisely, we can also allow $b_{L}\geq 1$ for the feature learning regime. However such initialization reduces to the case of $b_{L}=1$ in the next parameter update. Thus, we can summarize it as $b_{L}=1$. The scaling of $b_{L}=1$ implies that a smaller initialization is required compared to the standard parameterization (SP), which is PyTorch’s default, for sufficiently wide neural networks. It can also be immediately verified that we can set $a_{l}=0$ due to shift invariance without loss of generality. In Table 1, we summarize the $\mu$P from previous work and our results obtained in the following sections.